What nominal interest rate compounded continuously is equivalent to j52 = 14%?
Answer #1

Let the nominal interest rate = x

Hence, (1+x/52)52 = (1+0.14)

=> x = (1.141/52 - 1)*52 = 0.1312

Hence, x = 13.12%